ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 161, 2007

Acosta, M. D. and Sung Guen Kim Denseness of holomorphic functions attaining their numerical radii

373

Allen, P. Almost every 2-SAT function is unate  311
Anderson, G. W. Digit patterns and the formal additive group  125

Aschenbrenner, M. and H. Schoutens Lefschetz extensions, tight closure, and big Cohen-Macaulay algebras

221

Ceccherini-Silberstein, T. and M. Coornaert Injective linear cellular automata and sofic groups

1

Coornaert, M. See T. Ceccherini-Silberstein

Dilworth, S., E. Odell and B. Sari Lattice structures and spreading models

387

Fabińska, E. and M. Lassak Reduced bodies in normed planes  75
Greenberg, Matthew Lifting modular symbols of non-critical slope 141

Huang, Wen and Yingfei Yi A local variational principle and it applications to equilibrium states

29

Hug, D. and R. Schneider A stability result for a volume ratio  209

Kim, Sung Guen See M. D. Acosta

Klep, I. and M. Schweighofer A Nichtnegativstellensatz for polynomials in noncommuting variables

17

Kollár, J. Algebraic varieties over PAC fields  89

Laskowski, M. C. A simpler axiomatization of the Shelah-Spencer almost sure theories

157

Lassak, M. See E. Fabińska

Levcovitz , D., I. Vainsencher and F. Xavier Enumeration of cones over cubic scrolls

103

Narayanan, E. K., R. Rawat and S. K. Ray Approximation by Kfinite functions in L^{p} spaces

187

Odell, E. See S. Dilworth

Urbański, M. and A. Zdunik Instability of exponential Collet — Eckmann maps 347 Rawat, R. See E. Narayanan Ray, S. K. See E. Narayanan Sari, B. See S. Dilworth Schneider, R. See D. Hug Schoutens, H. See M. Aschenbrenner Schweighofer, M. See I. Klep Vainsencher, I. See D. Levcovitz Xavier, F. See D. Levcovitz Yi, Yingfei See Wen Huang Zdunik, A. See M. Urbański